OpenStudy (anonymous):

(y-4)/(y-6)-(y+1)/(y+6)+(y-30)/(y^2-36) simplify

jhonyy9 (jhonyy9):

(y-4)/(y-6)-(y+1)/(y+6)+(y-30)/(y*2-36)=((y+6)(y-4)-(y-6)(y+1)+(y-30))/((y-6)(y+6)) =((y*2+2y-24)-(y*2-5y-6)+(y-30))/((y-6)(y+6))=(8y-48)/((y-6)(y+6))= =8(y-6)/((y-6)(y+6))=8/(y+6)